find three consecutive even integers such that three times the middle integer is 4 more than the sum of the first and third.
I did 3(x+2)=x+x+4+4 and the answers I got were 6,8, and 10. is this correct?

Answer by sudhanshu_kmr(1152)


3(x+2)=x+(x+4)+4
=> 3x + 6 = 2x + 8
=> 3x - 2x = 8-6
=> x = 2
so, integers are 2, 4 and 6 .